OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = -2x - 8 and contains the point (4, 4).

OpenStudy (anonymous):

QUEG

OpenStudy (jhannybean):

a line that is perpendicular has a negative opposite slope of the original equation. So if your equation \(y=-2x-8\), \(m=-2\) The slope of the perpendicular line would be a negative reciprocal, so \(m=-(-\frac{1}{2})=2\) with this, and your points (4.4), use the point slope form. \[\large y-y_{1}=m(x-x_{1})\] plug in (4,4) as your (x,y) and m=2 as your slope. \[\large y-4=2(x-4)\] I'll let you simplify and solve it :)

OpenStudy (anonymous):

DINGO

OpenStudy (jhannybean):

Do you understand how to solve the rest, @lastkings ?

OpenStudy (anonymous):

is it y=2x+4?

OpenStudy (jhannybean):

You're almost right, but it's y= 2x -8+4